OpenStudy (anonymous):

What is the mean price per quart for each of the three stores? Store #1: $4.82 Store #2: $6.38 Store #3: $8.24 What is the weighted mean price per quart for all the stores together? Weighted mean: __________ How would you calculate the weighted mean? I found the means of the stores myself. That's easy enough. I'm just unsure of finding the weighted mean.

OpenStudy (austinl):

mean = average \[average = \frac{a_1 + a_2 + a_3 + ... + a_n}{n}\] \[= \frac{4.82 + 6.38 + 8.24}{3} = ?\]

OpenStudy (anonymous):

All I have to do is find the average of the means I already have? 6.48

OpenStudy (austinl):

Correct!

OpenStudy (anonymous):

You're awesome! Thanks!(:

OpenStudy (austinl):

No problemo.